# Fine-Tuning Masked Diffusion for Provable Self-Correction

### v2

@@ 1 Introduction
Concurrent work. [^1] This paragraph has been included from version two of this manuscript. Although entailed with different details, (meshchaninov2025guided; li2025riv) address the self-correction problem for MDM by training an additional module that takes $\hat{{\mathbf{x}}}_{0}$ as input, thereby belonging in the class of proxy-$\hat{{\mathbf{x}}}$ approaches in Section 2.2. The most relevant work, huang2025don, shares the same motivation and experimental setup but differs in several aspects, which we elaborate on in Appendix A.
@@ 4.2 Unconditional Text Generation
Figure 4: Assessing PRISM on unconditional text generation. PRISM (red) exhibits superior performance compared to baselines (ReMDM: green, ReMDM-conf: blue, Vanilla MDM: gray), particularly in the regime with fewer sampling steps ($N<1024$).
@@ 4.3 Appyling PRISM to 88B MDM
Results. We evaluate the resulting model in a zero-shot setting on MBPP (austin2021program), a challenging Python coding benchmark with $974$ problems. As shown in Table 1, PRISM outperforms ReMDM and ReMDM-conf, especially in the regime of $256$ sampling steps, by approximately $2.7\%$. We detail training and inference configurations in Appendix D.5D.6.
@@
PRISM 23.0% 30.5% 32.9%
ReMDM 20.2% 29.7% 33.2%
ReMDM-conf 20.3% 30.0% 32.7%
MDM 18.6% 28.2% 32.7%
Table 1: PRISM on LLaDA-8B-Instruct.
@@ 5 Conclusion
Discussion on experimental results. MDMs model per-position unmasking posteriors rather than the joint distribution over positions. Consequently, with few-step inference, they are prone to dependency errors across positions (liu2024discrete; xu2024energy), which creates a natural opportunity for our self-correction mechanism to detect erroneous predictions. We confirm this empirically: methods that incorporate remasking outperform vanilla MDM inference as the number of sampling steps decreases (Figure 4, Table 1). In particular, PRISM achieves larger gains than baselines by learning per-token quality. However, as the number of sampling steps increases, partially unmasked sequences contain fewer errors, leaving less headroom for the self-correction. For the scaling-up experiments in Section 4.3, although we fine-tune LLaDA-8B-Instruct on only 0.1M dataset pairs and evaluate it on MBPP, we expect that a more generalized model could be obtained by constructing a richer and more diverse dataset, given sufficient compute.
@@ Appendix A Related Works
Comparison to huang2025don. For clarity, we provide a comprehensive comparison with huang2025don. The overall training pipeline is largely similar but differs in several key aspects: (1) we provide theoretical guarantees for our training method (PRISM), and (2) we use a pretrained MDM to sample from $\mathbf{z}$} to $\mathbf{y}$}, whereas huang2025don construct $\mathbf{y}$} via a random corruption process. Nevertheless, their large-scale experiments on LLaDA-8B-Instruct cover a broader range of tasks and datasets, while our work focuses on relatively smaller-scale experiments.
@@ Appendix B Training algorithms and guarantees of PRISM
B.1 Marginal Equivalence of the Masking Process
@@
B.2 PRISM provably learns the per-token quality
@@
B.3 Extension on PRISM’s provable guarantee
@@ D.1 Omitted Results
Table 2: Evaluation result for PRISM-loop/PRISM/baselines. The best MAUVE scores are bolded.
@@ D.3 Ablation on Fine-tuning Hyperparameters-$(k,n_{y})$
We conducted an ablation study to analyze how the interplay between $k$ and $n_{y}$ affects the generation performance. We tested variants of $(k,n_{y})\in\{(4,8),(8,4),(16,2),(32,1)\}$, thus remaining the total number of token updates per batch constant ($k\times n_{y}=32$) for a fair comparison. All other configurations follow Table  3, except that we set nucleus $p=0.9$ to sample $\mathbf{y}$} from $\mathbf{z}$} (during the PRISM fine-tuning).
@@ D.4 Ablation on Fine-tuning Hyperparameters-Nucleus sampling
D.5 Experiment Configurations
@@
D.6 Additional experimental details on LLaDA

### v3

@@ 4.2 Unconditional Text Generation
Figure 4: Unconditional text generation performance. Metrics are evaluated at $\text{NFE}\in\{128,256,512,1024,2048,4096\}$; black dashed lines denote validation set references. PRISM (red) outperforms baselines (ReMDM: green, ReMDM-conf: blue, Vanilla MDM: gray), particularly at lower sampling steps ($N<1024$). Detailed numerical results are reported in Table 3.
Quantitative analysis. To further investigate PRISM’s behavior at a mechanistic level, we evaluate how well its per-token quality scores are calibrated to the desired likelihood. For partially masked sequences $\mathbf{x}_{t}$ encountered at inference, we collect per-token quality scores $g_{\phi}^{i}(\mathbf{x}_{t})$ over all clean positions $i$ and group these tokens into buckets according to the predicted scores. For each bucket, we then estimate an empirical oracle likelihood and query it by a pretrained model $\theta$; $f_{\theta}^{i}(\mathbf{x}_{t}^{i}\mid\mathbf{x}_{t}\oplus\mathbf{m}_{i})$ for each $i$. Interestingly, the resulting calibration curves (Figure 7) closely follow the queried likelihoods, indicating that $g_{\theta}^{i}(\mathbf{x}_{t})$ is a well-calibrated estimate of the target posterior, as desired.
@@ 4.3 Appyling PRISM to 88B MDM
Results. We evaluate the resulting model in a zero-shot setting on MBPP (austin2021program) and HumanEval (chen2021evaluating), a challenging Python coding benchmarks. As shown in Table 2, PRISM outperforms ReMDM and ReMDM-conf across different numbers of sampling steps. We detail training and inference configurations in Appendix D.6D.7.
@@
PRISM 28.0% 39.0% 42.7%
ReMDM 26.2% 34.8% 42.5%
ReMDM-conf 25.6% 34.1% 36.6%
MDM 25.6% 34.1% 36.6%
Table 1: HumanEval zero-shot performance.
Sampling steps 256 512 1024
PRISM 21.8% 29.1% 32.3%
ReMDM 19.7% 28.5% 32.9%
ReMDM-conf 20.1% 29.0% 32.5%
MDM 18.2% 27.8% 31.9%
Table 2: MBPP zero-shot performance.
@@ 5 Conclusion
Discussion on experimental results. MDMs model per-position unmasking posteriors rather than the joint distribution over positions. Consequently, with few-step inference, they are prone to dependency errors across positions (liu2024discrete; xu2024energy; kang2025parallelbench), which creates a natural opportunity for our self-correction mechanism to detect erroneous predictions. We confirm this empirically: methods that incorporate remasking outperform vanilla MDM inference as the number of sampling steps decreases (Figure 4, Table 2). In particular, PRISM achieves larger gains than baselines by learning per-token quality. However, as the number of sampling steps increases, partially unmasked sequences contain fewer errors, leaving less headroom for the self-correction. For the scaling-up experiments in Section 4.3, although we fine-tune LLaDA-8B-Instruct on only 0.1M dataset pairs, we expect that a more generalized model could be obtained by constructing a richer and more diverse dataset, given sufficient compute.
@@ Appendix B Training algorithms and guarantees of PRISM
B.1 PRISM fine-tuning algorithm
B.2 Marginal Equivalence of the Masking Process
@@
B.3 PRISM provably learns the per-token quality
@@
B.4 Extension on PRISM’s provable guarantee
@@ D.1 Omitted Results
Table 3: Evaluation result for PRISM-loop/PRISM/baselines. The best MAUVE scores are bolded. Baseline results marked with are taken from wang2025remasking.
@@ D.3 Ablation on Fine-tuning Hyperparameters-$(k,n_{y})$
We conducted an ablation study to analyze how the interplay between $k$ and $n_{y}$ affects the generation performance. We tested variants of $(k,n_{y})\in\{(4,8),(8,4),(16,2),(32,1)\}$, thus remaining the total number of token updates per batch constant ($k\times n_{y}=32$) for a fair comparison. All other configurations follow Table  4, except that we set nucleus $p=0.9$ to sample $\mathbf{y}$} from $\mathbf{z}$} (during the PRISM fine-tuning).
@@ D.4 Ablation on Fine-tuning Hyperparameters-Nucleus sampling
D.5 Quantitative analysis of PRISM
We provide the calibration analysis result of PRISM in Figure 7.
Figure 7: Calibration of PRISM per-token quality on OpenWebText during generation without remasking with $256$ steps. For each timestep, tokens are binned by predicted quality $g_{\theta}^{i}(\mathbf{x}_{t})$, and for each bin we plot the empirical probability that token $i$ matches the ground truth, estimated via $f_{\theta}(\cdot\mid\mathbf{x}_{t}\oplus\mathbf{m}_{i})$, against the bin mean.
D.6 Experiment Configurations
@@
D.7 Additional experimental details on LLaDA